# Automatically Identifying Local and Global Circuits with Linear Computation Graphs

### v1

@@ Appendix C Hierachical Attribution Algorithm
In this section, we introduce the detailed implementation of Hierachical Attribution algorithm, to obtain a subgraph $G^{\prime}$ from original computational graph $G$ with threshold $\tau$, as shown in Algorithm 1.
@@
Afterwards, we can compute the contribution of $G^{\prime}$ by adding up the attribution score of all its leaf nodes.
@@ Appendix D Equality of Output Activation and Leaf Nodes Attribution
The activation of output node $t$ is also a linear function of all leaf nodes by iteratively applying $a_{u}=\sum_{v\to u\in E}k_{v,u}a_{v}$, which is
$a_{t}$ $=\sum_{\deg_{\text{in}}(v)=0}a_{v}\cdot\nabla_{a_{t}}a_{v}$ (7)
@@ Appendix E Additional Explanation of IOI Circuit
  • "John", "and" and "Mary" features simply imply current token as "John", "and", and "Mary";
  • Entity Indicator features are activated on prepositions or transitive verbs, indicating that its next token is likely to be an entity.
  • "John" Preceding features collect information from the previous token, and imply its previous token as "John";
  • "And" Preceding features collect information from the previous token, and imply its previous token as "and";
  • Consecutive Entity features are a mixture of "Mary" features and "And" Preceding features, and imply the current token as the [B] part of an [A] and [B] pattern, where [A] and [B] serve as entities.
@@
  • "And"-Connected Entities Preceding features collect information from several previous tokens (mainly the token "and"), and imply there’s an [A] and [B] pattern before this token.
@@
  • Centered Entity Association features take advantage of the structural information from "And"-Connected Entities Induction features, and decide to retrieve the entity lying before "and", by attending to Centered Entity features in Name Mover Heads. This behavior is not completely symmetrical to that with Consecutive Entity features, since Centered Entity features do not know about the token "and" after it. However, this behavior is still reasonable since if there’s another Centered Entity before IO, then this entity can be another correct answer.

### v2

@@ Appendix C Hierachical Attribution Algorithm
In this section, we introduce the detailed implementation of the Hierarchical Attribution algorithm to obtain a subgraph $G^{\prime}$ from the original computational graph $G$ with threshold $\tau$, as shown in Algorithm 1.
@@
Afterwards, we can compute $G^{\prime}$’s contribution by adding up the attribution scores of all its leaf nodes.
@@ Appendix D Equality of Output Activation and Leaf Nodes Attribution
For any activated node $u$ (i.e., $a_{u}>0$), it holds that
$a_{u}$ $=\operatorname{ReLU}\left(\sum_{v\to u\in E}k_{v,u}a_{v}\right)$ (10)
$=\sum_{v\to u\in E}k_{v,u}a_{v}$
$=\sum_{v\to u\in E,a_{t}>0}k_{v,u}a_{v}$
By iteratively applying Eq. 10, we can obtain
$a_{t}$ $=\sum_{\deg_{\text{in}}(v)=0,a_{v}>0}a_{v}\cdot\nabla_{a_{t}}a_{v}$ (11)
$=\sum_{\deg_{\text{in}}(v)=0}a_{v}\cdot\nabla_{a_{t}}a_{v}$
@@ Appendix E Additional Explanation of IOI Circuit
  • "John", "and" and "Mary" features simply imply the current token as "John", "and", and "Mary";
  • Entity Indicator features are activated on prepositions or transitive verbs, indicating that its next token will likely be an entity.
  • "John" Preceding features collect information from the previous token and imply its previous token as "John";
  • "And" Preceding features collect information from the previous token and imply its previous token as "and";
  • Consecutive Entity features are a mixture of "Mary" features and "And" Preceding features imply the current token as the [B] part of an [A] and [B] pattern, where [A] and [B] serve as entities.
@@
  • "And"-Connected Entities Preceding features collect information from several previous tokens (mainly the token "and") and imply there’s an [A] and [B] pattern before this token.
@@
  • Centered Entity Association features take advantage of the structural information from "And"-Connected Entities Induction features and decide to retrieve the entity lying before "and", by attending to Centered Entity features in Name Mover Heads. This behavior is not completely symmetrical to that with Consecutive Entity features since Centered Entity features do not know about the token "and" after it. However, this behavior is still reasonable since if there’s another Centered Entity before IO, then this entity can be another correct answer.